OpenStudy (anonymous):

Water freezes at 32 F or 0 C. Use temperature conversion formulas to show that these two measures are indeed the same temperature.

OpenStudy (anonymous):

@godgavemeyou @Hope_nicole and anyone else that can help :)

OpenStudy (anonymous):

look up a temp conversion formula

OpenStudy (anonymous):

It's given that water freezes at 0 C or 32 F We know that temp in Fahrenheit (F) is related to temp in Celsius(C) by this relation F=95×C+32 Here C=0, so F=0+32=32 So the two temperatures are the same

OpenStudy (anonymous):

@Hope_nicole I tried looking it up, couldn't find one. or one I understand. lol i'm not good at math at all. @godgavemeyou Thanks... Could you dummer it down please? Im really dum when it comes to math.

OpenStudy (anonymous):

she gave you the equation. F=95*C+32. that is the conversion from celcius to Farenheit.

OpenStudy (amistre64):

i developed the conversion formulas when i was a kid .... then i learned about them later. we have 2 points to create a linear relationship with C F C F (0,32) and (100,212) using the 2 points we can form a slope: 212-32 ------- = 180/100 = 18/10 = 9/5 100 - 0 pluging this slope and a known point into the point slope form of a line we get: F-32 = 9/5 (C-0) F = 9/5 C + 32 , which we can solve for C as well C = 5/9 (F - 32)

OpenStudy (anonymous):

@Hope_nicole Thank you, sorry for the trouble. I am realllllyyyy stupid when it comes to math. -_- lol

OpenStudy (anonymous):

@amistre64 Thank you. but I already got an answer that i understand.

OpenStudy (amistre64):

thats ok :)

OpenStudy (anonymous):

Could you help me with one more question ?? @amistre64

OpenStudy (amistre64):

maybe, ive only got so many smarts about me in a day ... i might be running low :)

OpenStudy (anonymous):

lol